May rides her bike the same distance that Leah walks. May rides her bike 10 km per hour faster than Leah walks. If it takes May 1 hour and Leah 3 hours to travel that distance, how far does each travel. What is the formula? Please help I am stuck.
Answer by Nate(3500)

Rate * Time = Distance
Distance = Distance
p is going to be defined as Leah's rate
May~>
rate: 10 + p
time = 1
Leah~>
rate: p
time: 3
(10 + p)(1) = (p)(3)
10 + p = 3p
10 = 2p
5 = p
Leah's rate is 5km per hour. It takes her three hours to walk the distance.
Rate * Time = Distance
5 * 3 = 15km